OpenStudy (anonymous):

find the exact value of tan 112.50

OpenStudy (dumbcow):

\[\tan 112.5 = \frac{\sin 122.5}{\cos 112.5} = \frac{\sin (90+\frac{45}{2})}{\cos (90+\frac{45}{2})} = \frac{\cos (\frac{45}{2})}{-\sin (\frac{45}{2})}\] use half angle identity

OpenStudy (anonymous):

this is the whole equation?

OpenStudy (dumbcow):

look up the half angle identities for sin/cos plug in values for sin(45/2) and cos(45/2) simplify

OpenStudy (anonymous):

my answer is 2!!is this correct?

OpenStudy (anonymous):

hi

OpenStudy (dumbcow):

nope not quite http://www.wolframalpha.com/input/?i=tan%28112.5%29&a=TrigRD_D

OpenStudy (anonymous):

find the exact value of sin 112.5 also

OpenStudy (dumbcow):

what? well you are doing that in the process of getting tan(112.5) lol

OpenStudy (dumbcow):

if you show your work i can assist in where you went wrong